NetflixAndNo (netflixandno):

What is the value of c so that x^2 + 17x + c is a perfect square trinomial?

OpenStudy (mathstudent55):

We are dealing with a perfect square trinomial with a leading coefficient of 1. That means that the coefficient of the x^2 term is 1. \((x+a)^2 = (x + a)(x + a) = x^2 + ax + ax + a^2 = x^2 + 2ax + a^2\) After we square the binomial ax + a, we get \(x^2 + 2ax + a^2\). Notice that the middle term is 2ax. The last term is a^2. If you divide 2a by 2, you get a. Then square a to get a^2. That means if you take half of the middle coefficient and square it, you get the last term of a perfect square trinomial that has leading coefficient 1.

NetflixAndNo (netflixandno):

Thank you!

OpenStudy (mathstudent55):

You're welcome.